# Sorting inversion sequences

### v1

@@ 3.1 Restricting pop stack depth
When extending the notion of the depth of a stack to words, we could continue to use the definition in terms of the number of elements allowed in the stack as introduced by Murray [10]. However, given the repetition that distinguishes words from permutations, we will instead use the following definition that is also consistent with the original definition applied to permutations.
@@
The depth of a stack (or pop stack) will be the number of entries with distinct values allowed in the stack (or pop stack) at any stage.
For example, a stack of depth three under this definition would still only allow three entries of a permutation to be in the stack at one time. However, this same restricted stack could allow arbitrarily many entries of a word in the stack, provided that no more than three distinct letters were represented at any time. We begin by considering pop stacks of limited depth.
@@
A word is sortable by a pop stack of depth $k$ if and only if avoids $120,201,1010$ and also $k(k-1)\cdots 10$.
@@
Conversely, if $w$ is not pop stack sortable by a pop stack of depth $k$ and avoids $120,201,1010$, then the problem is a result of the limitation of the pop stack’s capacity. Hence $w$ must contain a $k(k-1)\cdots 10$ pattern. ∎
@@
An inversion sequence is sortable by a pop stack of depth $k$ if and only if avoids $120,201,1010$ and also $k(k-1)\cdots 10$.
In the case of depth one, the number of sortable inversion sequences is the number of weakly increasing inversion sequences which are known to be enumerated by the Catalan numbers due to Martinez and Savage [24]. The proof shows these inversion sequences to correspond to the $213$ avoiding permutations enumerated by Simion and Schmidt [26].
@@
For a pop stack of depth two, the sortable inversion sequences would need to avoid $120,201,210,1010$. The first few terms of the enumeration sequence are: $1,1,2,6,23,100,471,2349$. We can define the generating tree for these inversion sequences (using the algorithm introduced in [16]) as having a root $a_{1}$ and satisfying the following rules
$a_{m}$ $\rightsquigarrow a_{m+1},b_{m,1},\ldots,b_{m,m},\quad m\geq 1,$
$c_{m}$ $\rightsquigarrow c_{m+1},a_{m+2},b_{m+1,1},\ldots,b_{m+1,m+1},\quad m\geq 1,$
where $a_{m}=0^{m}$, $b_{m,j}=0^{m}j$ and $c_{m}=0^{m}10$. However, it is hard to get an explicit formula by translating this generating tree to system of equations and is even more difficult for larger values of $k$.
@@
For all $n,k\in\mathbb{Z}^{+}$, with $k<n$
$|WD_{k}(n,a)|=1+\sum_{j=1}^{k}\binom{n-1}{j}\binom{a-1}{j-1}$ (1)
Note we assume $\binom{n}{k}=0$ if $k>n$. The result follows from choosing $j$ locations for a descent, then $j-1$ values for the weakly decreasing sequence, and one for the sequence with no descents.
We say a word is a layered weakly decreasing word of depth at most $k$ and length $n$ if it can be written as a sequence of subwords $\{w_{1},w_{2},...\}$ where each $w_{i}\in WD_{k}(m_{i},a_{i})$ with $\sum_{i}m_{i}=n$, and for each $i$, $a_{i}\geq 0$, and $min(w_{i+1})\geq max(w_{i})$ (i.e. the largest (first) value of each word is bounded by the smallest value of the succeeding word). Further, we consider the words of this type which are also inversion sequences, i.e. $max(w_{i})\leq\sum_{j=1}^{i-1}m_{j}$. Let $LDI_{k}(n)$ be the set of all such layered depth at most $k$ inversion sequences of length $n$.
@@
Given an element $\{w_{1},w_{2},...w_{t}\}\in LDI_{k}(n)$ as above with each value of $w_{i+1}$ bounded below by the largest value in $w_{i}$, subtracting $a_{i}$ from every value in $w_{i+1}$ yields an element of $WD_{k}(m_{i},a_{i+1}-a_{i})$.
@@
$|LDI_{k}(n)|=\sum_{m_{1}+m_{2}+...m_{t}=n,m_{i}\geq 0}\sum_{a,1,a_{2},...a_{t}}^{*}|WD(m_{i},a_{i})|,$ (2)
where the $\sum^{*}$ means to only sum on values of ${0\leq a_{j}\leq\bigg{(}\sum_{h=1}^{j-1}m_{h}-a_{h}\bigg{)}-1}$.
Since the first word in an element of $LDI_{k}(n)$ must be a string of zeros, we let $LDI_{k}(n,m_{1})$ denote the elements of $LDI_{k}(n)$ that begin with $m_{1}$ zeros followed by a non-zero value. Further, let $LDI_{k}(n,m_{1},a_{2})$ is the subset of these layered decreasing inversion sequences which begin with $m_{1}$ zeros, and first non-zero value $a_{2}$ (and $a_{2}\leq m_{1}$). We now decompose our sequence of words as $\{\overline{0}_{m_{1}},w_{2},\hat{w}\}$ where $\hat{w}=\{w_{3},...w_{t}\}$. Now $w_{2}\in WD_{k}(m_{2},a_{2})$ for some $0\leq m_{2}\leq n-m_{1}$, $a_{2}>0$. If we subtract $a_{2}$ from all the elements in $\hat{w}$ the result is in $LDI(n-m_{2}-m_{1},m_{1}+m_{2}-a_{2},a)$ for some $a$ with $a\leq m_{1}+m_{2}-a_{2}$. Note that the number of starting zeros guarantees that adding $a_{2}$ to all of these entries will still yield an element of $LDI_{k}(n)$.
Hence we get the formula,
@@
$LDI_{k}(n,m_{1})=\sum_{m_{2}=0}^{n-m_{1}}\sum_{a_{2}\leq m_{1}}WD_{k}(m_{2},a_{2})\sum_{a\leq m_{1}+m_{2}-a_{2}}LDI_{k}(n-m_{1}-m_{2},m_{1}+m_{2}-a_{2},a).$ (3)
Both of the above formulas are rather difficult to use for computation. However they are presented to give some insight into directly enumerating these inversion sequences.
@@
A word is sortable by a stack of depth $k$ if and only if avoids $120$ and also $k(k-1)\cdots 10$.
@@
The number of words of length $n$ on an alphabet $[k]$ sortable by a stack of depth $1$, i.e. avoid $10$ is $\binom{n+k-1}{k-1}$.
@@
Finally, we note Kotsireas, Yıldırım, and the first author [16] give a functional equation for the inversion sequences of length $n$ avoiding $120,210$ and thereby sortable by a stack of depth $2$.
@@ 3.2 Generalizing pop stacks
Let an $(r,1)$-stack allow the push operation to push an entry to any of the top $r$ “value positions” in the stack where consecutive entries of the same value in the stack are grouped together. Note that the same definition could be extended to the broader notion of an $(r,s)$ stack where we allow the pop operation to be defined in terms of these same value positions.
@@
[…]
Figure 3: A $(3,1)$-stack
@@
$w=w_{1}w_{2}\cdots w_{n}$ is a word of length $[n]$.
@@
$i=1$ is the index of the next entry in the input
$j=0$ is the counter for the number of copies of the smallest value in the $(2,1)$-Pop Stack of Depth $2$
@@
By way of contradiction let $w$ be a sortable word that is not sorted by our algorithm on a $(2,1)$-pop stack of depth 2. Specifically, $S(w)$ must contain a descent, say $w_{j}w_{k}$. Because Algorithm 3 never allows a descent in the pop stack, $w_{j}$ must have been the last entry popped from one pop operation and $w_{k}$ must have been the first entry popped from the next pop operation. The entries $w_{j}$ and $w_{k}$ must have appeared in the same relative order (forming an inversion) in $w$. The only way to remove this inversion would be for both entries to be in the pop stack at the same time.
@@
The enumeration problem for this $(2,1)$-pop stack of depth $2$ can be considered for permutations, words, and specifically inversion sequences. The permutations of length $n$ avoiding $231,312,321$, i.e. $120,201,210$, are known as the layered permutations where each layer can have size one or two. Note that layered permutations are exactly the permutations that avoid $231,312$. Further, the restriction of avoiding $321$ forces each layer (consecutive decreasing subsequence) to have size at most two. These restricted layered permutations are enumerated by the Fibonacci numbers as shown by Simion and Schmidt [26].
@@
[…]
Burstein [5, Theorem 5.1], enumerated the words of length $n$ over alphabet $[k]$ avoiding $120,201,210$.
@@
(Burstein) The generating function for the number of words of length $n$ over alphabet $[k]$ sortable by a $(2,1)$-pop stack of depth $2$ is given by the coefficient of $x^{n}y^{k}$ in $F(x,y)$ where
@@
The case for the enumeration of sortable inversion sequences by a $(2,1)$-pop stack of depth $2$ (in the form of avoidance of $120,201,210$ or otherwise) does not appear in the literature prior to this work, but can be solved using generating trees as demonstrated in subsection 4.1. Specifically, the generating function for these sortable inversion sequences is given in Theorem 4.3.
@@ 4.   The generating tree method
To enumerate the inversion sequences of length $n$ avoiding $120,201,210$, i.e. the inversion sequences sortable by a $(2,1)$-pop stack of depth $2$, we use the generating tree method shown by Kotsireas, Yıldırım, and the first author [16].
Let ${\mathbf{A}}=\cup_{n=0}^{\infty}{\mathbf{I}}_{n}(120,201,210)$ be the set of all nonempty inversion sequences that avoid $120$, $201$, and $210$. We construct a pattern-avoidance tree ${\mathbf{T}}$ for the class of pattern-avoiding inversion sequences ${\mathbf{A}}$ as follows.
@@
Now, we relabel the vertices of the tree ${\mathbf{T}}$ as follows. Define ${\mathbf{T}}(e)$ to be the subtree consisting of the inversion sequence $e$ as the root and its descendants in ${\mathbf{T}}$. We say that $e$ is equivalent to $e^{\prime}$, denoted by $e\sim e^{\prime}$, if and only if ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$ (in the sense of plane trees). Let ${\mathbf{T}}^{\prime}$ be the same tree ${\mathbf{T}}$ where we replace each node $e$ by the first node $e^{\prime}\in{\mathbf{T}}$ from top to bottom and from left to right in ${\mathbf{T}}$ such that ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$.
@@
Also, the children of $b_{m,j}\in{\mathbf{T}}^{\prime}$ are $b_{m,j}0,b_{m,j}1,\ldots,b_{m,j}(m+1)$. Notice ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+2-j,1})$ with $k=0,1,\ldots,j-1$. Clearly, when $j=1$, any inversion sequence $0^{m}10\pi^{\prime}$ in ${\mathbf{I}}_{n}(120,201,210)$ if any only if $0^{m+1}1\pi^{\prime}$ in ${\mathbf{I}}_{n}(120,201,210)$. Thus, we can assume that $11\leq j\leq k-$ 1. To prove the claim for such $j$, let $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$. Since $j>k$ and $\pi$ avoids both $201$ and $210$, we see that there are no letters in $\pi^{\prime}$ belonging to the set $\{0,\ldots,k-1,k+1,\ldots,j-1\}$. Thus, by mapping $\pi$ to $0^{m+2-k}1\pi^{\prime\prime}$, where $\pi^{\prime\prime}$ is obtained from $\pi^{\prime}$ by mapping each letter $x$ of $\pi^{\prime}$ to $x+1-j$ with $x\neq k$, and mapping the letter $k$ to 00. Hence, we see that $\pi\in{\mathbf{I}}_{n}(120,201,210)$ if and only if $0^{m+2-j}1\pi^{\prime\prime}\in I_{n+2-j}(120,201,210)$.
@@
Define $A_{m}(x)$ (respectively, $B_{m,j}(x)$) to be the generating function for the number of nodes at level $n\geq 1$ for the subtree of ${\mathbf{T}}(B;a_{m})$ (respectively, ${\mathbf{T}}(B;b_{m,j})$), where the root stays at level 00. Thus, by Lemma 4.2, we have
@@
where the last equation (8) is the translating of (5) with $j=1$.
@@
By using expressions of $A(0),C(0)$, we obtain an explicit formula for $A(v)$ from Equation (11). We can then obtain an explicit formula for $C(v)$ from Equation (10). Then we can use that formula to get an explicit formula for $B(v,1)$ from Equation (9). Finally, we have the information we need to obtain an explicit formula for $B(v,u)$ from Equation (7). We omit the presentations of the expressions $A(v),C(v),B(v,1),B(v,u)$ because they are very lengthy. However, these expressions are satisfy the system of Equations (6)-(8). Hence, we can state the following result.

### v5

@@ 3.1 Restricting pop stack depth
When extending the notion of the depth of a stack to words, we could continue to use the definition in terms of the number of elements allowed in the stack as introduced by Elder [14]. This could be more practical in some applications. However, given the repetition that distinguishes words from permutations and considering the storage in terms of distinct elements to remember, it also makes sense to consider another definition that is still consistent with the original definition when applied to permutations. The content of the stack could be stored as a limited number (the depth) of lists of positions in the stack where these positions are counted from the bottom. We consider the latter notion as given in the following definition.
@@
The depth of a stack (or pop stack) is the number of entries with distinct values allowed in the stack (or pop stack) at any stage.
For example, a stack of depth three under this definition would still only allow three entries of a permutation to be in the stack at one time. However, this same restricted stack could allow arbitrarily many entries of a word in the stack, provided that no more than three distinct letters were represented at any time. Using this definition, we begin by characterizing words and specifically inversion sequences sortable by pop stacks of limited depth.
@@
A word $w$ is sortable by a pop stack of depth $k$ if and only if $w$ avoids $120,201,1010$ and also $k(k-1)\cdots 10$.
@@
Conversely, if $w$ is not pop stack sortable by a pop stack of depth $k$ and avoids $120,201,1010$, then the problem is a result of the limitation of the pop stack’s capacity. Hence $w$ must contain the $k(k-1)\cdots 10$ pattern. ∎
@@
Inversion sequences sortable by a pop stack of depth $k$ are exactly those in
${\mathbf{I}}(120,201,1010,k(k-1)\cdots 10)$.
In the case of depth one, the number of sortable inversion sequences is the number of weakly increasing inversion sequences which are known to be enumerated by the Catalan numbers due to Martinez and Savage [28]. The proof shows these inversion sequences to correspond to the $213$ avoiding permutations, which were enumerated by Simion and Schmidt [31].
@@
For a pop stack of depth two, the sortable inversion sequences would need to avoid $120,201,210,1010$. The first few terms of the enumeration sequence are: $1,1,2,6,23,100,471,2349$. We define the succession rules for the generating tree for these inversion sequences (using the algorithm introduced in [20]) in Section 4.3. However, it is hard to get an explicit formula by translating this generating tree to a system of equations and it is even more difficult for larger values of $k$.
@@
For all $n,k\in\mathbb{Z}^{+}$ with $k<n$, we have
$|WD_{k}(n,a)|=1+\sum_{j=1}^{k}\binom{n-1}{j}\binom{a}{j}.$ (1)
Note that we assume $\binom{n}{k}=0$ if $k>n$. The result follows from choosing $j$ locations for a descent, then $j$ values for the weakly decreasing sequence, and one for the sequence with no descents.
We say a word $w$ of length $n$ is a layered weakly decreasing word of depth at most $k$ if $w$ can be split into a sequence of subwords (in the factor sense) $w_{1},w_{2},...$ where each $w_{i}\in WD_{k}(m_{i},a_{i})$ with $\sum_{i}m_{i}=n$, and for each $i$, $a_{i}\geq 0$, and $min(w_{i+1})\geq max(w_{i})$ (i.e. the largest (first) value of each word is bounded by the smallest value of the succeeding word). Further, we consider the words of this type which are also inversion sequences, i.e. $max(w_{i})\leq\sum_{j=1}^{i-1}m_{j}$. Let $LDI_{k}(n)$ be the set of all such layered depth at most $k$ inversion sequences of length $n$.
@@
Let $w_{1}w_{2}...w_{t}\in LDI_{k}(n)$, where $w_{1},w_{2},...,w_{t}$ are the subwords as defined above with each value of $w_{i+1}$ bounded below by the largest value in $w_{i}$. Subtracting $a_{i}$ from every value in $w_{i+1}$ yields an element of $WD_{k}(m_{i},a_{i+1}-a_{i})$.
@@
${}|LDI_{k}(n)|=\sum_{\begin{subarray}{c}m_{1}+m_{2}+...m_{t}=n,\\ m_{i}\geq 0\end{subarray}}\quad\sum_{a_{1},a_{2},...a_{t}}^{*}|WD_{k}(m_{i},a_{i})|,$ (2)
where $a_{1}=0$, and the $\sum^{*}$ means for all $j>1$ to only sum on values of $a_{j}$ such that ${0\leq a_{j}\leq\bigg{(}\sum_{h=1}^{j-1}m_{h}-a_{h}\bigg{)}}$.
Since the first word in an element of $LDI_{k}(n)$ must be a string of zeros, we let $LDI_{k}(n,m_{1})$ denote the elements of $LDI_{k}(n)$ that begin with $m_{1}$ zeros followed by a non-zero value. Further, let $LDI_{k}(n,m_{1},a_{2})$ be the subset of these layered decreasing inversion sequences which begin with $m_{1}$ zeros, and have first non-zero value $a_{2}$ (and $a_{2}\leq m_{1}$). We now decompose our sequence of words as $\{\overline{0}_{m_{1}},w_{2},\hat{w}\}$ where $\hat{w}=\{w_{3},...w_{t}\}$. Now $w_{2}\in WD_{k}(m_{2},a_{2})$ for some $0\leq m_{2}\leq n-m_{1}$, $a_{2}>0$. If we subtract $a_{2}$ from all the elements in $\hat{w}$ the result is in $LDI(n-m_{2}-m_{1},m_{1}+m_{2}-a_{2},a)$ for some $a$ with $a\leq m_{1}+m_{2}-a_{2}$. Note that the number of starting zeros guarantees that adding $a_{2}$ to all of these entries will still yield an element of $LDI_{k}(n)$.
Hence we get the following formula.
@@
${}|LDI_{k}(n,m_{1})|=\sum_{m_{2}=0}^{n-m_{1}}\sum_{a_{2}\leq m_{1}}|WD_{k}(m_{2},a_{2})|\sum_{a\leq m_{1}+m_{2}-a_{2}}|LDI_{k}(n-m_{1}-m_{2},m_{1}+m_{2}-a_{2},a)|.$ (3)
Equations 2 and 3 are rather difficult to use for computation. However they are presented to give some insight into directly enumerating these inversion sequences.
@@
A word $w$ is sortable by a stack of depth $k$ if and only if $w$ avoids $120$ and also $k(k-1)\cdots 10$.
@@
The number of words of length $n$ on an alphabet $[k]$ sortable by a stack of depth $1$, i.e. avoiding $10$ is $\binom{n+k-1}{k-1}$.
@@
Finally, we note that Kotsireas, Yıldırım, and the first author [20] give a functional equation for the inversion sequences of length $n$ avoiding $120,210$ and thereby sortable by a stack of depth $2$.
@@ 3.2 Generalizing pop stacks
Let an $(r,1)$-stack allow the push operation to push an entry to any of the top $r$ “value positions” in the stack where consecutive entries of the same value in the stack are grouped together.
Note that the above definition could be extended to the broader notion of an $(r,s)$ stack where we allow the pop operation to be defined in terms of these same value positions.
@@
230230230230666666664444$\square$3333$\square$0000$\square$303030306666666644443333$\square$2222$\square$0000$\square$0066666666444433333333$\square$2222$\square$0000$\square$66666666444433333333$\square$2222$\square$000000$\square$
Figure 3: A $(3,1)$-stack shown at several stages.
@@
$w=w_{1}w_{2}\cdots w_{n}$ is a word of length $n$.
@@
$i=1$ is the index of the next entry in the input.
$s=0$ is the counter for the number of copies of the smallest value in the $(2,1)$-Pop Stack of Depth $2$.
@@
By way of contradiction, let $w$ be a sortable word that is not sorted by our algorithm on a $(2,1)$-pop stack of depth 2. Specifically, $S(w)$ must contain a descent, say $w_{j}w_{k}$. Because Algorithm 3 never allows a descent in the pop stack, $w_{j}$ must have been the last entry popped from one pop operation and $w_{k}$ must have been the first entry popped from the next pop operation. The entries $w_{j}$ and $w_{k}$ must have appeared in the same relative order (forming an inversion) in $w$. The only way to remove this inversion would be for both entries to be in the pop stack at the same time.
@@
The enumeration problem for this $(2,1)$-pop stack of depth $2$ can be considered for permutations, words, and specifically inversion sequences. The permutations of length $n$ avoiding $231,312,321$, i.e. $120,201,210$, are known as the layered permutations where each layer can have size one or two. Note that layered permutations are exactly the permutations that avoid $231,312$. Furthermore, the restriction of avoiding $321$ forces each layer (consecutive decreasing subsequence) to have size at most two. These restricted layered permutations are enumerated by the Fibonacci numbers as shown by Simion and Schmidt [31]. As a consequence of their result and Theorem 3.23 we have the following corollary.
@@
(Simion and Schmidt) The number of permutations of length $n$ sortable by a $(2,1)$-pop stack of depth $2$ is the $(n+1)^{\text{st}}$ Fibonacci number $F_{n+1}$.
Burstein [6, Theorem 5.1] enumerated the words of length $n$ over an alphabet $[k]$ avoiding $120,201,210$.
@@
(Burstein) The generating function for the number of words of length $n$ over an alphabet $[k]$ sortable by a $(2,1)$-pop stack of depth $2$ is given by the coefficient of $x^{n}y^{k}$ in $F(x,y)$ where
@@
The case for the enumeration of sortable inversion sequences by a $(2,1)$-pop stack of depth $2$ (in the form of avoidance of $120,201,210$ or otherwise) does not appear in the literature prior to this work, but can be solved using generating trees as demonstrated in Section 4.1. Specifically, the generating function for these sortable inversion sequences is given in Theorem 4.3.
@@ 4.   The generating tree method
To enumerate the inversion sequences of length $n$ avoiding $120,201,210$, i.e. the inversion sequences sortable by a $(2,1)$-pop stack of depth $2$, we use the generating tree method shown by Kotsireas, Yıldırım, and the first author [20]. See the recent paper of Pantone [29] for another nice description of using succession rules to enumerate pattern avoiding inversion sequences.
Let ${\mathbf{A}}=\cup_{n=0}^{\infty}{\mathbf{I}}_{n}(120,201,210)$. We construct a pattern-avoidance tree ${\mathbf{T}}$ for the class of pattern-avoiding inversion sequences ${\mathbf{A}}$ as follows.
@@
Now, we relabel the vertices of the tree ${\mathbf{T}}$ as follows. Define ${\mathbf{T}}(e)$ to be the subtree consisting of the inversion sequence $e$ as the root and its descendants in ${\mathbf{T}}$. We say that $e$ is equivalent to $e^{\prime}$, denoted by $e\sim e^{\prime}$, if and only if ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$ (in the sense of plane trees). Define an order on the nodes of any tree to be from top to bottom, and within a level from left to right. Denote by ${\mathbf{T}}^{\prime}$ the tree ${\mathbf{T}}$ where we have replaced, in order, each node $e\in T$ by the first node $e^{\prime}\in{\mathbf{T}}$ (from top to bottom and within a level from left to right) in ${\mathbf{T}}$ such that ${\mathbf{T}}(e)\cong{\mathbf{T}}(e^{\prime})$ assuming that such a node $e^{\prime}$ exists.
@@
Note that exponents are used in two ways below based on context. Specifically, when describing succession rules, an exponent of $j$ on a child indicates the number of copies of that child the parent has. For example, in Lemma 4.2, $b_{m,j}$ has $j$ children of the form $b_{m+2-j,1}$. However, when describing the individual nodes that make up the tree, the exponent refers to copies of the letter in the word (in particular, inversion sequence). Again, referring to Lemma 4.2, we have $a_{m}$ is the word made up of exactly $m$ 00s.
@@
Also, the children of $b_{m,j}\in{\mathbf{T}}^{\prime}$ are $b_{m,j}0,b_{m,j}1,\ldots,b_{m,j}(m+1)$. Notice that ${\mathbf{T}}(b_{m,j}k)\cong{\mathbf{T}}(b_{m+2-j,1})$ with $k=0,1,\ldots,j-1$. To show this, we map any inversion sequence $\pi=0^{m}jk\pi^{\prime}\in{\mathbf{I}}_{n}(120,201,210)$ to $0^{m+2-j}1\pi^{\prime\prime}$, where $\pi^{\prime\prime}$ is obtained from $\pi^{\prime}$ (there are no letters in $\pi^{\prime}$ belonging to the set $\{0,\ldots,k-1,k+1,\ldots,j-1\}$) by replacing each letter $x\geq j$ by $x+1-j$ and replacing the letter $k$ by 00. Hence, we see that $\pi\in{\mathbf{I}}_{n}(120,201,210)$ if and only if $0^{m+2-j}1\pi^{\prime\prime}\in I_{n+2-j}(120,201,210)$.
@@
Define $A_{m}(x)$ (respectively, $B_{m,j}(x)$) to be the generating function for the number of nodes at level $n\geq 1$ for the subtree of ${\mathbf{T}}(B;a_{m})$ (respectively, ${\mathbf{T}}(B;b_{m,j})$), where the root stays at level $1$. Thus, by Lemma 4.2, we have
@@
where Equation (8) is the translation of (5) with $j=1$.
@@
Using the expressions of $A(0),C(0)$, we obtain an explicit formula for $A(v)$ from Equation (11). We can then obtain an explicit formula for $C(v)$ from Equation (10). Then we can use that formula to get an explicit formula for $B(v,1)$ from Equation (9). Finally, we have the information we need to obtain an explicit formula for $B(v,u)$ from Equation (7). We omit the presentations of the expressions $A(v),C(v),B(v,1),B(v,u)$ because they are very lengthy. However, these expressions satisfy the system of Equations (6)-(8). Hence, we can state the following result.